OpenStudy (anonymous):

Need help with Trig Identities (1-tan(x))/(1-cot(x))=-tan (x)

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

it didnt show up?

OpenStudy (anonymous):

its (1-tan(x))/(1-cot(x)0=-tan(x)

OpenStudy (anonymous):

sorry there shouldnt b a zero after cot (x) it was suppose to b a bracket

OpenStudy (anonymous):

So the fraction can be rewritten as: $$ \frac{1-tan(x)}{1-cot(x)} = \frac{1-\frac{sin(x)}{cos(x)}}{1- \frac{cos(x)}{sin(x)}} $$

OpenStudy (anonymous):

Now multiply top and bottom by $$sin(x) \cdot cos(x)$$

OpenStudy (anonymous):

$$\frac{1-\frac{sin(x)}{cos(x)}}{1- \frac{cos(x)}{sin(x)}} \cdot \frac{sin(x)cos(x)}{sin(x)cos(x)}$$

OpenStudy (anonymous):

cant draw it to me cuz i cant see how ur doing that

OpenStudy (anonymous):

y do u have to multiply it sin x * cos x

OpenStudy (anonymous):

Because now the equation becomes: $$ \frac{sin(x)cos(x) - sin^2(x)}{sin(x)cos(x) - cos^2(x)} $$

OpenStudy (anonymous):

oh ok i see how u got there

OpenStudy (anonymous):

now what do u do

OpenStudy (anonymous):

The equation can be written as $$\frac{sin(x)(cos(x) - sin(x))}{cos(x)(sin(x) - cos(x))}$$

OpenStudy (anonymous):

what happened to the squares? how did u get rid of them?

OpenStudy (anonymous):

I pulled out a $$ sin(x) $$ from the top, and a $$cos(x)$$ from the bottom Notice that if you distribute them, you will get back $$\frac{sin(x)cos(x) - sin^2(x)}{sin(x)cos(x) - cos^2(x)}$$

OpenStudy (anonymous):

w8 then shouldnt it be \[\frac{ 1*\cos(x)-\sin(x) }{ \sin(x)*1-\cos(x) }\]

OpenStudy (anonymous):

if u take factor out sin (x) and cos (x)

OpenStudy (anonymous):

Well yes, it's the same thing without the 1's

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You *could*, but then that would defeat the purpose of multiplying the fraction by $$sin(x)cos(x)$$ in the first place

OpenStudy (anonymous):

So, now we can rewrite the fraction as : $$\frac{sin(x)}{cos(x)} \cdot \frac{(cos(x) - sin(x))}{(sin(x) - cos(x))}$$

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

And we know that $$\frac{sin(x)}{cos(x)}$$ is $$tan(x)$$

OpenStudy (anonymous):

So the fraction is: $$tan(x) \cdot \frac{cos(x) - sin(x)}{sin(x) - cos(x)}$$

OpenStudy (anonymous):

oh ok then u can cancel out cos x-sin x/sin x-cos x which will give u -1

OpenStudy (anonymous):

oh w8 no

OpenStudy (anonymous):

Well, you can't cancel, because when you have a plus or minus, you can't cancel. You have to treat the top and bottom as one whole entity.

OpenStudy (anonymous):

If you give me a minute, I can look up the identity required, but I know you can turn the second fraction into -1.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Uggh. This is going to bug me. I know it's negative one, but I forget why.

OpenStudy (anonymous):

i got it up to cosx(sinx)-1+sinx(cos x)

OpenStudy (anonymous):

but i think that wrong

OpenStudy (anonymous):

Oh... duh. Yeah. Just like if you had: $$\frac{(4-7)}{(7-4)} = \frac{-3}{3} = -1$$ This works the same way.

OpenStudy (anonymous):

So the top and bottom will always be the same value, but will have different signs. Therefore, is will always be negative one.

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

i dont got how to do get that with this

OpenStudy (anonymous):

wat.

OpenStudy (anonymous):

how do u get -1 with (cos x-sin x)/(sin x-cos x)

OpenStudy (anonymous):

is there an identity for cosx(sinx)?

OpenStudy (anonymous):

There is no trig identity or anything. Say cos(x)-sin(x) = some value m Then we can say sin(x)-cos(x) will be -m. This is just like my example above: if $$(7-4) = m$$ then $$m=3$$ We also know that $$(4-7) = -3$$ So you can see that when you flip the numbers in a subtraction problem, you will just get the negative result. $$\frac{(4-7)}{(7-4)} = \frac{-m}{m} = \frac{-3}{3} = -1$$

OpenStudy (anonymous):

Sorry if that's not clear enough. It's not anything complicated, I think you're overthinking it.

OpenStudy (anonymous):

So: If (a - b) = m. We can conclude that (b - a) will be -m.

OpenStudy (anonymous):

so u cant get -1 with out acutally solving the expression u just have to know its -1?

OpenStudy (anonymous):

Yeah... one of the unfortunate parts of Calc & Trig.

OpenStudy (anonymous):

well that sucks but ok thanks

OpenStudy (anonymous):

Haha no problem. Hope it helped.

OpenStudy (anonymous):

took 2 pages of writing for this question lol